OpenStudy (anonymous):

FInd the vertex, focus, and directix of the parabola. x=2y^2

OpenStudy (amistre64):

a parabola has the expression: y^2 = 4ax ....or some similar topology to it..

OpenStudy (amistre64):

if we rearrange your expression we get: (1/2)x = y^2 which gets us looking close to normal again

OpenStudy (amistre64):

4a = 1/2 a = 1/8 ; the distance from the vertex to the directix os 1/8 ; and the distance from the vertex to the focus is 1/8

OpenStudy (amistre64):

the parabola opens to the left, and it is centered onthe xaxis; so.... vertex = (0,0) focus = (0,1/8) and directix is the line x= -1/8

OpenStudy (anonymous):

thanks =)

OpenStudy (amistre64):

youre welcome.... its always good to review the basics ;)